Find the differential equation of the family of curves y=ae^x+be^-x
, where
A and B are arbitrary constants.

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
dy/dx =ae^x- be^-x
d^2y/ dx^2 = ae^x+ be^-x
 
hence d^2y/dx^2 = y
 
hope it helps
 
Regards
Arun (askIITians forum expert)